Can 89/13, 151/11 or 223/23 be simplified? Even though they are reduced to the lowest terms, but can they be simplified? Which ones can or none of the above. Thank you

Answer by AnlytcPhil(1806)

Can 89/13, 151/11 or 223/23 be simplified?  Even though they are reduced to the
lowest terms, but can they be simplified?  Which ones can or none of the
above.  Thank you

They can only be changed to improper fractions, if you can call that
"being simplified":

 89/13 = 6 11/13

151/11 = 13 8/11

223/23 = 9 16/23

Edwin
AnlytcPhil@aol.com